A 3-year bond with a 10% coupon rate paid annually and a $1,000 face value sells at a nominal yield
to maturity 8% (APR). What is the price of the bond?

N:

I/Y:

PV:

PMT:

FV:

Mode:

Excel Formula:

Bond Proceeds:

Quoted Bond Price:
Answer #1

Price of the Bond = 100* 2.577 +1000*0.7938= 1051.50

N: 3*1=3

I/Y: 0.08/1 =0.08

PV of interest payment: 100*2.577=257.7

PMT: 1000*.10

FV:1000

Bond Proceeds:1000*0.7938=793.80

Quoted Bond Price:1051.50